OpenStudy (marcelie):

help please... im kind of struggling with my algebra so i have this as my answer...

OpenStudy (marcelie):

\[u ^{-1}\] so then would this be my answer ? \[\frac{u ^{-1} }{ -1 }\]

OpenStudy (marcelie):

so does the exponent of -1 go on the denominator ?

OpenStudy (saadsword):

u^(-1)=1/u^(1)= 1/u The negative exponent goes in the denominator an turns positive

OpenStudy (marcelie):

oh okay so then would it be - 1 /u ?

OpenStudy (s4sensitiveandshy):

the confusing part i guess is that you're comparing this with a anti-derivative power rule

OpenStudy (saadsword):

no 1/u

OpenStudy (saadsword):

only exponents change their signs

OpenStudy (marcelie):

can you give me an example lol

OpenStudy (s4sensitiveandshy):

\[u^{-1}\] can be written as \[\frac{ u^{-1} }{ 1 }\] to convert negative to positive exponent flip the fraction ( as mentioned above) so it would be \[\frac{ 1 }{ u^1 }\]

OpenStudy (marcelie):

hmm i think im confusing myself.. let me post my work... one sec

OpenStudy (s4sensitiveandshy):

\[\int\limits_{ }^{ } u^{-2}\] for this one add one to the power and then divide by the new power \[\frac{ u^{-2+1}}{ -2+1 } = \frac{ u^{-1} }{ -1 } = -\frac{1}{u}\]

OpenStudy (marcelie):

its the last step in my paper where i did u sub

OpenStudy (s4sensitiveandshy):

i just checked that last part (u sub) looks good to me

OpenStudy (marcelie):

ah okay so thats where i struggled for the u^-2 so then the - got factored out right

OpenStudy (s4sensitiveandshy):

well don't forget the integral sign right there \[\int\limits_{ }^{ } u^{-2} d \theta\] what would be the next step ?

OpenStudy (marcelie):